An iron wire weighs 2.225g. It’s dissolved in an acid which oxidizes to Fe2+ and made up of 250cm3. A 25cm3 of sample required is 31cm3 of a 0.0185mol L-1 solution of potassium dichromate in a titration.

  1. Calculate the number of moles of potassium dichromate used in the titration?
  2. Determine number of moles of iron present in the 250cm3 iron wire solution?
  3. Calculate mass of iron present in the iron wire?
  4. Calculate percentage of iron present in the iron wire?
1
Expert's answer
2021-08-09T07:47:28-0400

Molar Mass of iron = 55.485

Molar Mass of potassium dichromate =294.185

= 18.5 × 31 = 573.5

= 573.5/294.185

= 1.95moles